# Continuous Invariance Learning

### v1

@@ Appendix G Proof for Lemma 1
$\operatorname*{arg\,min}_{h\in\mathcal{H}}\mathbb{E}_{(\boldsymbol{x},t)}(h(\Phi(\boldsymbol{x}))-\boldsymbol{t})^{2}$
$=$ $\operatorname*{arg\,min}_{h\in\mathcal{H}}\mathbb{E}_{(\Phi(\boldsymbol{x}),t)}(h(\Phi(\boldsymbol{x}))-\boldsymbol{t})^{2}$
$=$ $\operatorname*{arg\,min}_{h\in\mathcal{H}}\mathbb{E}_{\Phi(\boldsymbol{x})}\mathbb{E}_{\boldsymbol{t}\sim\mathbb{P}(\boldsymbol{t}|\Phi(\boldsymbol{x}))}(h(\Phi(\boldsymbol{x}))-\boldsymbol{t})^{2}.$
@@
$\mathbb{E}_{\boldsymbol{t}\sim\mathbb{P}(\boldsymbol{t}|\Phi(\boldsymbol{x}))}(h(\Phi(\boldsymbol{x}))-\boldsymbol{t})^{2}$
$=$ $h(\Phi(\boldsymbol{x}))^{2}-2h(\Phi(\boldsymbol{x}))^{\top}\mathbb{E}[\boldsymbol{t}|\Phi(\boldsymbol{x})]+\mathbb{E}[\boldsymbol{t}^{2}|\Phi(\boldsymbol{x})],$
we then know the minimum is achieved at $h(\Phi(\boldsymbol{x}))=\mathbb{E}[\boldsymbol{t}|\Phi(\boldsymbol{x})]=h^{*}(\Phi(\boldsymbol{x}))$ by solving this quadratic problem. The minimum loss achieved by $h^{*}(\cdot)$ is
$\mathbb{E}_{\Phi(\boldsymbol{x})}\left[\mathbb{E}[\boldsymbol{t}^{2}|\Phi(\boldsymbol{x})]-(\mathbb{E}[\boldsymbol{t}|\Phi(\boldsymbol{x})])^{2}\right]=\mathbb{E}_{\Phi(\boldsymbol{x})}[\mathbb{V}[\boldsymbol{t}|\Phi(\boldsymbol{x})]],$
We can prove the result for $g^{*}(\cdot)$ in a similar way, and the minimum loss achieved by $g^{*}(\cdot)$ is $\mathbb{E}_{\Phi(\boldsymbol{x}),\boldsymbol{y}}[\mathbb{V}[\boldsymbol{t}|\Phi(\boldsymbol{x}),\boldsymbol{y}]]$. ∎
@@ Appendix H Proof for Theorem 2
The penalty term of Eqn (3) is
$\min_{h\in\mathcal{H}}\max_{g\in\mathcal{G}}\mathbb{E}_{\boldsymbol{x},\boldsymbol{y},\boldsymbol{t}}\left[(h(\Phi(\boldsymbol{x}))-\boldsymbol{t})^{2}_{2}-(g(\Phi(\boldsymbol{x}),\boldsymbol{y})-\boldsymbol{t})^{2}_{2}\right]$
$=$ $\mathbb{E}_{\Phi(\boldsymbol{x})}[\mathbb{V}[\boldsymbol{t}|\Phi(\boldsymbol{x})]]-\mathbb{E}_{\Phi(\boldsymbol{x}),\boldsymbol{y}}[\mathbb{V}[\boldsymbol{t}|\Phi(\boldsymbol{x}),\boldsymbol{y}]]$
$=$ $\mathbb{E}_{\Phi(\boldsymbol{x}),\boldsymbol{y}}[\mathbb{E}[\boldsymbol{t}|\Phi(\boldsymbol{x}),\boldsymbol{y}]^{2}]-\mathbb{E}_{\Phi(\boldsymbol{x})}[\mathbb{E}[\boldsymbol{t}|\Phi(\boldsymbol{x})]^{2}]$
$=$ $\mathbb{E}_{\Phi(\boldsymbol{x}),\boldsymbol{y}}[\mathbb{E}[\boldsymbol{t}|\Phi(\boldsymbol{x}),\boldsymbol{y}]^{2}]-\mathbb{E}_{\Phi(\boldsymbol{x})}[(\mathbb{E}_{\boldsymbol{y}}\mathbb{E}[\boldsymbol{t}|\Phi(\boldsymbol{x}),\boldsymbol{y}])^{2}]$
@@
The last inequality is due to Jensen’s inequality and the convexity of the quadratic function. The inequality is achieved only when $\mathbb{E}[\boldsymbol{t}|\Phi(\boldsymbol{x})]=\mathbb{E}[\boldsymbol{t}|\Phi(\boldsymbol{x}),\boldsymbol{y}],\forall\boldsymbol{y}\in\mathcal{Y}$. ∎
Appendix I Proof for Theorem 2
@@
Assumption 4 (Li and Liu [2021]).
[…]
@@
[Li and Liu [2021]] Denoting $(w,\Phi,h)$ as $\theta$ and $Q(w,\Phi,h,g)$ as $Q(\theta,g)$ for short, $Q(\theta,g)$ is $\mu$-strongly convex in $\theta$, i.e.,
@@
The convex-concave assumption for the minimax problem is popular in existing literature Li and Liu [2021], Lei et al. [2021], Farnia and Ozdaglar [2021], Zhang et al. [2021], simply because non-convex-non-concave minimax problems are extremely hard to analyze due to their non-unique saddle points. When $w$ and $h$ (the classifier for $\boldsymbol{y}$ and regressor for $\boldsymbol{t}$) are linear, it is easy to verify that $Q$ is convex in them. Furthermore, recent theoretical studies show that the overparameterized neural networks (NN) behave like convex systems and training large NN is likely to converge to the global optimum Jacot et al. [2018], Mei et al. [2018].
Assumption 6 (Lipschitz continuity [Li and Liu, 2021]).
@@
Assume we solve (3.1) by SGDA as introduced in Algorithm 1 and obtain $(\epsilon_{1},\epsilon_{2})$ solution $-^(\hat{\theta},\hat{h})$. Further, suppose Assumption 5 holds and SGDA is $\epsilon$-argument-stable as described in Assumption 4, then for any $\delta>0$, fix $\eta>0$, we have with probability at least $1-\delta$
@@
In this section, we show the spurious relationship $p_{s}(\boldsymbol{t})$ varies across the domains as shown in Figure 4.
@@
Appendix K Additional Results for the Experiments 4.1.1
The result across different split number is shown in Table 9
@@
None ERM 0.2533 (0.0201) 0.3618 (0.0306)
Discrete 2 IRMv1 0.5311 (0.0192) 0.6616 (0.0429)
4 0.5539 (0.0687) 0.6991 (0.0240)
5 0.3935(0.1170) 0.5598(0.0391)
8 0.4790 (0.0244) 0.5903 (0.0532)
16 0.3066 (0.0445) 0.5690 (0.0498)
50 0.2732(0.0097) 0.4273(0.1147)
100 0.2590(0.0285) 0.4128(0.1366)
2 REx 0.3969 (0.0456) 0.5976 (0.0122)
4 0.5580 (0.1249) 0.6405 (0.1129)
5 0.5077(0.1169) 0.6515(0.0577)
8 0.4646 (0.7500) 0.6815 (0.0551)
16 0.4405 (0.0953) 0.6778 (0.0660)
50 0.4273(0.1147) 0.6118(0.0278)
100 0.4187(0.0048) 0.6333(0.0505)
Continuous CIL 0.6095 (0.0659) 0.7625 (0.0499)
Table 6: Comparison on the synthetic Logit data. Metric is accuracy and standard deviations in parenthesis are calculated on three independent runs. Split Num stands for the number of domains we manually create by equally splitting the raw continuous domains. We try settings where the spurious correlation $p_{s}(\boldsymbol{t})$ changes by linear and Sine functions, respectively.
Appendix L Additional Experiment Results for 4.1.2
Details on continous CMNIST. The original CMNIST dataset consists of two domains with varying spurious correlation values: 0.9 in one domain and 0.8 in the other. To simulate a continuous problem, we randomly assign time indices 1-512 to samples in the first domain and indices 513-1024 to the second domain. The spurious correlation only changes at time index 513, as shown in the Figure 5 in the Page 16 of the Appendix file. In this case, each of the 1024 domains comprises approximately 50 samples. Based on the results in Figure 2 in the main part of the manuscript, REx and IRMv1 display testing accuracies close to random guessing in this scenario. Similar constructions are made for 4, 8, …, 512 domains.
In this section, we provide the complete experiment results in Table 7 and Table 8 for IRMv1, REx, GroupDRO and IIBNet with other number of splits (“split num") on the the continuous CMNIST dataset. The settings are the same as in Section 4.1.2, where Table 3show the best results for each method with one corresponding “split num".
@@
Discrete 2 IRMv1 0.4974 (0.0062) 0.5001 (0.0034)
4 0.5102 (0.0086) 0.4972 (0.0086)
8 0.5145 (0.0353) 0.4946 (0.0094)
16 0.5131 (0.0228) 0.4971 (0.0090)
100 0.5021 (0.0196) 0.4935(0.0005)
500 0.4956 (0.0074) 0.4702(0.0272)
1000 0.4967 (0.0033) 0.4870(0.0265)
2 REx 0.8309 (0.0040) 0.4550 (0.0112)
4 0.8275 (0.0033) 0.4714 (0.0222)
8 0.8205 (0.0067) 0.4931 (0.0255)
16 0.8232 (0.0045) 0.4786 (0.0154)
100 0.8054 (0.0078) 0.4912(0.0020)
500 0.7921 (0.0045) 0.4255(0.0105)
1000 0.7898 (0.0032) 0.4187(0.0048)
Table 7: Accuracy on the continuous CMNIST with $p_{s}(\boldsymbol{t})$ as a linear function with different split number in (2, 4, 8, 16).
@@
Discrete 2 IRMv1 0.4947 (0.0021) 0.4944 (0.0057)
4 0.4977 (0.0078) 0.4996 (0.0055)
8 0.4960 (0.0065) 0.4985 (0.0026)
16 0.4974 (0.0062) 0.5001 (0.0034)
100 0.4962 (0.0058) 0.4983(0.0063)
500 0.5028 (0.0087) 0.4623(0.1011)
1000 0.5043 (0.0123) 0.4963(0.1506)
2 REx 0.8295 (0.0047) 0.5417 (0.0137)
4 0.8193 (0.0091) 0.5497 (0.0171)
8 0.8159 (0.0089) 0.5496 (0.0210)
16 0.8160(0.0069) 0.5407 (0.0179))
100 0.8094 (0.0078) 0.4866(0.0079)
500 0.8148 (0.0067) 0.4314(0.0097))
1000 0.7997 (0.0079) 0.4224(0.0074)
Table 8: Accuracy on the continuous CMNIST with $p_{s}(\boldsymbol{t})$ as a Sine function with different split number in (2, 4, 8, 16).
Results. Table 7 and 8 report the training and testing accuracy of different domain splitting schemes (2, 4, 8, 16) for both IRMv1 and REx on the continuous CMNIST dataset with two $p_{s}(\boldsymbol{t})$ settings. Under linear $p_{s}(\boldsymbol{t})$, REx performs the best with split number 8 by improving around 4% over the worst one. However, the results are fairly close under sine $p_{s}(\boldsymbol{t})$. On the other hand, IRMv1 with all testing accuracy around 0.5 does not seem to be able to extract useful features for the task with either linear $p_{s}(\boldsymbol{t})$ or sine $p_{s}(\boldsymbol{t})$.
@@
Discrete 2 IRMv1 0.7268(0.036) 0.7152(0.0135)
5 0.7346(0.0141) 0.6841(0.0113)
10 0.7297(0.0188) 0.6728(0.0164)
25 0.7442(0.0210) 0.5425(0.0175)
50 0.7540(0.0099) 0.5209(0.0205)
2 REx 0.7097(0.0006) 0.7443 (0.0132)
5 0.7130(0.0117) 0.7442(0.0087)
10 0.7046(0.0025) 0.7320(0.0165)
25 0.7101(0.0130) 0.7296(0.0024)
50 0.6882(0.0092) 0.7290(0.0046)
Continuous CIL 0.6095 (0.0659) 0.7625 (0.0499)
Table 9: Comparison on the real world dataset: HousePrice and Insuran . Metric is accuracy and standard deviations in parenthesis are calculated on three independent runs. Split Num stands for the number of domains we manually create by equally splitting the raw continuous domains.
@@
In this section, we tried different penalty set up in Eqn (3) on the auto-scaling dataset to validate the robustness of CIL. The approximating functions $h(\Phi(\boldsymbol{x})),g(\Phi(\boldsymbol{x}),\boldsymbol{y})$ are implemented by 2-Layer MLPs. We evaluate the performance changes by increasing either the hidden dimension of the MLPs (Table 10) or penalty weight $\lambda$ (Table 11).
Penalty Dimension Train Test
32 0.8437 (0.0595) 0.6424 (0.0547)
64 0.8417 (0.0258) 0.7040 (0.0118)
128 0.8232 (0.0538) 0.6936 (0.0213)
256 0.8557 (0.0267) 0.6812 (0.0225)
512 0.8528 (0.0227) 0.6877 (0.0160)
Table 10: Training and Testing accuracy on the auto-scaling dataset across different MLP setups for $h(\Phi(\boldsymbol{x})),g(\Phi(\boldsymbol{x}),\boldsymbol{y})$ in Eqn (3). Standard deviation in brackets is calculated with 3 independent runs. Other settings are kept the same as in Section O
@@
100 0.8691 (0.0131) 0.6760 (0.0126)
1000 0.8447 (0.0257) 0.7001 (0.0154)
10000 0.8417 (0.0258) 0.7040 (0.0118)
100000 0.8395 (0.0273) 0.7042 (0.0116)
1000000 0.8395 (0.0273) 0.7035 (0.0124)
Table 11: Training and Testing accuracy on the auto-scaling dataset across different penalty weights $\lambda$ in Eqn (3). Standard deviation in brackets is calculated with 3 independent runs. Other settings are kept the same as in Section O
Results. The training and testing accuracy of different MLP setup (hidden dimension) for $h(\Phi(\boldsymbol{x})),g(\Phi(\boldsymbol{x}),\boldsymbol{y})$ under CIL on the auto-scaling dataset are shown in Table 10. CIL performs the best when the hidden dimension is 64, and the performance is stable with even higher dimensions. However, the testing accuracy drops to 64% when the dimension is reduced to 32, as the model is probably too simple to be able to extract enough invariant features. But it is still better than ERM shown in Table 4.
@@
  • LR: learning rate of the classification model $\Phi(\boldsymbol{x}))$, e.g. 1e-3.
  • OLR: learning rate of the penalty model $h(\Phi(\boldsymbol{x})),g(\Phi(\boldsymbol{x}),\boldsymbol{y})$, e.g. 0.001
@@
We show the parameter values used for each dataset in the following tables.

### v2

@@ Appendix G Proof for Lemma 1
$\operatorname*{arg\,min}_{h\in\mathcal{H}}\mathbb{E}_{(\bm{x},t)}(h(\Phi(\bm{x}))-\bm{t})^{2}$
$=$ $\operatorname*{arg\,min}_{h\in\mathcal{H}}\mathbb{E}_{(\Phi(\bm{x}),t)}(h(\Phi(\bm{x}))-\bm{t})^{2}$
$=$ $\operatorname*{arg\,min}_{h\in\mathcal{H}}\mathbb{E}_{\Phi(\bm{x})}\mathbb{E}_{\bm{t}\sim\mathbb{P}(\bm{t}|\Phi(\bm{x}))}(h(\Phi(\bm{x}))-\bm{t})^{2}.$
@@
$\mathbb{E}_{\bm{t}\sim\mathbb{P}(\bm{t}|\Phi(\bm{x}))}(h(\Phi(\bm{x}))-\bm{t})^{2}$
$=$ $h(\Phi(\bm{x}))^{2}-2h(\Phi(\bm{x}))^{\top}\mathbb{E}[\bm{t}|\Phi(\bm{x})]+\mathbb{E}[\bm{t}^{2}|\Phi(\bm{x})],$
we then know the minimum is achieved at $h(\Phi(\bm{x}))=\mathbb{E}[\bm{t}|\Phi(\bm{x})]=h^{*}(\Phi(\bm{x}))$ by solving this quadratic problem. The minimum loss achieved by $h^{*}(\cdot)$ is
$\mathbb{E}_{\Phi(\bm{x})}\left[\mathbb{E}[\bm{t}^{2}|\Phi(\bm{x})]-(\mathbb{E}[\bm{t}|\Phi(\bm{x})])^{2}\right]=\mathbb{E}_{\Phi(\bm{x})}[\mathbb{V}[\bm{t}|\Phi(\bm{x})]],$
We can prove the result for $g^{*}(\cdot)$ in a similar way, and the minimum loss achieved by $g^{*}(\cdot)$ is $\mathbb{E}_{\Phi(\bm{x}),\bm{y}}[\mathbb{V}[\bm{t}|\Phi(\bm{x}),\bm{y}]]$. ∎
@@ Appendix H Proof for Theorem 2
The penalty term of Eqn equation 3 is
$\min_{h\in\mathcal{H}}\max_{g\in\mathcal{G}}\mathbb{E}_{\bm{x},\bm{y},\bm{t}}\left[(h(\Phi(\bm{x}))-\bm{t})^{2}_{2}-(g(\Phi(\bm{x}),\bm{y})-\bm{t})^{2}_{2}\right]$
$=$ $\mathbb{E}_{\Phi(\bm{x})}[\mathbb{V}[\bm{t}|\Phi(\bm{x})]]-\mathbb{E}_{\Phi(\bm{x}),\bm{y}}[\mathbb{V}[\bm{t}|\Phi(\bm{x}),\bm{y}]]$
$=$ $\mathbb{E}_{\Phi(\bm{x}),\bm{y}}[\mathbb{E}[\bm{t}|\Phi(\bm{x}),\bm{y}]^{2}]-\mathbb{E}_{\Phi(\bm{x})}[\mathbb{E}[\bm{t}|\Phi(\bm{x})]^{2}]$
$=$ $\mathbb{E}_{\Phi(\bm{x}),\bm{y}}[\mathbb{E}[\bm{t}|\Phi(\bm{x}),\bm{y}]^{2}]-\mathbb{E}_{\Phi(\bm{x})}[(\mathbb{E}_{\bm{y}}\mathbb{E}[\bm{t}|\Phi(\bm{x}),\bm{y}])^{2}]$
@@
The last inequality is due to Jensen’s inequality and the convexity of the quadratic function. The inequality is achieved only when $\mathbb{E}[\bm{t}|\Phi(\bm{x})]=\mathbb{E}[\bm{t}|\Phi(\bm{x}),\bm{y}],\forall\bm{y}\in\mathcal{Y}$. ∎
Appendix I Proof for Proposition 2
@@
Assumption 4 ((Li & Liu, 2021)).
Let $\bar{\mathcal{D}}^{\text{tr}}$ be the dataset generated by replacing one data point in the training dataset $\mathcal{D}^{\text{tr}}$ with another data point drawn independently from the training distribution. We assume SGDA is $\epsilon$-argument-stable if for any $\mathcal{D}^{\text{tr}}$ such that
$\|\theta_{\text{SGDA}}(\bar{\mathcal{D}}^{\text{tr}})-\theta_{\text{SGDA}}(\mathcal{D}^{\text{tr}})\|\leq\epsilon,\quad\|g_{\text{SGDA}}(\bar{\mathcal{D}}^{\text{tr}})-g_{\text{SGDA}}(\mathcal{D}^{\text{tr}})\|\leq\epsilon$ (5)
@@
[(Li & Liu, 2021)] Denoting $(w,\Phi,h)$ as $\theta$ and $Q(w,\Phi,h,g)$ as $Q(\theta,g)$ for short, $Q(\theta,g)$ is $\mu$-strongly convex in $\theta$, i.e.,
@@
The convex-concave assumption for the minimax problem is popular in existing literature (Li & Liu, 2021; Lei et al., 2021; Farnia & Ozdaglar, 2021; Zhang et al., 2021), simply because non-convex-non-concave minimax problems are extremely hard to analyze due to their non-unique saddle points. When $w$ and $h$ (the classifier for $\bm{y}$ and regressor for $\bm{t}$) are linear, it is easy to verify that $Q$ is convex in them. Furthermore, recent theoretical studies show that the overparameterized neural networks (NN) behave like convex systems and training large NN is likely to converge to the global optimum (Jacot et al., 2018; Mei et al., 2018).
Assumption 6 (Lipschitz continuity (Li & Liu, 2021)).
@@
Assume we solve equation 3.1 by SGDA as introduced in Algorithm 1 and obtain $(\epsilon_{1},\epsilon_{2})$ solution $(\hat{\theta},\hat{h})$. Further, suppose Assumption 5 holds and SGDA is $\epsilon$-argument-stable as described in Assumption 4, then for any $\delta>0$, fix $\eta>0$, we have with probability at least $1-\delta$
@@
In this section, we show the spurious relationship $p_{s}(\bm{t})$ varies across the domains as shown in Figure 6.
@@
Appendix K Logit Dataset Experiment
Figure 7: Spurious correlation of logit dataset
Setting. We generate the first synthetic dataset with the invariant feature $\bm{x}_{v}\in\mathbb{R}^{2}$, spurious feature $\bm{x}_{s}\in\mathbb{R}^{20}$ and target $\bm{y}\in\{0,1\}$. The continuous domains $\bm{t}\in[0,100]$. The conditional distribution of $\bm{x}_{s}$ given $\bm{y}$ varies along $\bm{t}$ as follows:
$\bm{y}=$$\begin{cases}0,\mbox{ w.p. }0.5,\\ 1,\mbox{ w.p. }0.5,\end{cases}\bm{x}_{v}\sim$$\begin{cases}\mathcal{N}(y,\sigma^{2}),\mbox{ w.p. }p_{v},\\ \mathcal{N}(-y,\sigma^{2}),\mbox{ w.p. }1-p_{v},\end{cases}\bm{x}_{s}\sim$$\begin{cases}\mathcal{N}(y,\sigma^{2}),\mbox{ w.p. }p_{s}(\bm{t}),\\ \mathcal{N}(-y,\sigma^{2}),\mbox{ w.p. }1-p_{s}(\bm{t}),\end{cases}$
where $p_{v}$ and $p_{s}(\bm{t})$ are the probabilities of the feature’s agreement with label $\bm{y}$.
Notably, $p_{s}(\bm{t})$ varies among domains while $p_{v}$ stays invariant. The observed feature $\bm{x}$ is a concatenation of $\bm{x}_{v}$ and $\bm{x}_{s}$, i.e., $\bm{x}:=[\bm{x}_{v},\bm{x}_{s}]$. We generate 2000 samples as the training dataset $\{(\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{t}_{i})\}_{i=1}^{2000}$. Our goal is to learn a model $f(\bm{x})$ to predict $\bm{y}$ that solely relies on $\bm{x}_{v}$. We set $p_{v}$
@@
None ERM 25.33(2.01) 36.18(3.06)
Discrete 4 IRMv1 55.39(6.87) 69.91(2.40)
8 REx 46.46(7.50) 68.15(5.51)
2 GroupDRO 46.68(1.80) 60.95(2.23)
16 IIBNet 51.50(18.60) 49.68(21.87)
Continuous IRMv1 25.90(2.85) 41.28(3.66)
REx 42.73(11.47) 63.33(5.05)
Diversify 53.39(2.96) 53.30(1.93)
CIL 60.95(6.59) 76.25(4.99)
Table 4: Comparison on the synthetic Logit data. The metrics including accuracy and standard deviations in parenthesis are calculated on three independent runs. The environment type “Discrete” indicates that we manually create by equally splitting the raw continuous domains. The environment type “Continuous” indicates we use the original continuous domain index. “Split Num” stands for the number of domains we manually create and we report the best performance among spilt $\{2,4,8,16\}$.
to be 0.9 in all domains. $p_{s}(\bm{t})$ varies across the domains as in Figure 7 .
We can see that $\bm{x}_{s}$ exhibits a high but unstable correlation with $\bm{y}$. Since existing methods need discrete domains, we equally divide the continuous domain into different numbers of discrete domains, i.e, $\{2,4,8,16\}$.
Results. Table 4 shows the test accuracy of each method. ERM performs the worst and the large gap implies that ERM models heavily depend on spurious features. IRMv1 improves by 25% on average compared to ERM. CIL outperforms ERM by over 30% on average. CIL also improves by 5-7% over all existing methods based on discrete domains on the best manual discrete domain partition, which indicates that CIL can learn invariant features more effectively. The trend of how the performance of IRMv1 and REx changes with the split number is shown in Table 5
Env. Type Split Num Method Linear $p_{s}$ Sine $p_{s}$
None ERM 25.33(2.01) 36.18(3.06)
Discrete 2 IRMv1 53.11(1.92) 66.16(4.29)
4 55.39(6.87) 69.91(2.40)
5 39.35(11.70) 55.98(3.91)
8 47.90(2.44) 59.03(5.32)
16 30.66(4.45) 56.90(4.98)
50 27.32(0.97) 42.73(11.47)
100 25.90(2.85) 41.28(13.66)
2 REx 39.69(4.56) 59.76(1.22)
4 55.80(12.49) 64.05(11.29)
5 50.77(11.69) 65.15(5.77)
8 46.46(0.75) 68.15(5.51)
16 44.05(9.53) 67.78(6.60)
50 42.73(11.47) 61.18(2.78)
100 41.87(0.48) 63.33(5.05)
Continuous CIL 60.95(6.59) 76.25(4.99)
Table 5: Comparison on the synthetic Logit data. The metric is accuracy and standard deviations in parenthesis are calculated on three independent runs. Split Num stands for the number of domains we manually create by equally splitting the raw continuous domains. We try settings where the spurious correlation $p_{s}(\bm{t})$ changes by linear and Sine functions, respectively.
Appendix L Additional Experiment Results for 4.1.1
Details on continous CMNIST. The original CMNIST dataset consists of two domains with varying spurious correlation values: 0.9 in one domain and 0.8 in the other. To simulate a continuous problem, we randomly assign time indices 1-512 to samples in the first domain and indices 513-1024 to the second domain. The spurious correlation only changes at time index 513, as shown in Figure 5 on Page 16 of the Appendix file. In this case, each of the 1024 domains comprises approximately 50 samples. Based on the results in Figure 2 in the main part of the manuscript, REx and IRMv1 display testing accuracies close to random guessing in this scenario. Similar constructions are made for 4, 8, …, 512 domains.
In this section, we provide the complete experiment results in Table 6 and Table 7 for IRMv1, REx, GroupDRO, and IIBNet with other numbers of splits (“split num") on the continuous CMNIST dataset. The settings are the same as in Section 4.1.1, where Table 2show the best results for each method with one corresponding “split num".
@@
Discrete 2 IRMv1 49.74(0.62) 50.01(0.34)
4 51.02(0.86) 49.72(0.86)
8 51.45(3.53) 49.46(0.94)
16 51.31(2.28) 49.71(0.90)
100 50.21(1.96) 49.35(0.05)
500 49.56(0.74) 47.02(2.72)
1000 49.67(0.33) 48.70(2.65)
2 REx 83.09(0.40) 45.50(1.12)
4 82.75(0.33) 47.14(2.22)
8 82.05(0.67) 49.31(2.55)
16 82.32(0.45) 47.86(1.54)
100 80.54(0.78) 49.12(0.20)
500 79.21(0.45) 42.55(1.05)
1000 78.98(0.32) 41.87(0.48)
Table 6: Accuracy on the continuous CMNIST with $p_{s}(\bm{t})$ as a linear function with different split number in (2, 4, 8, 16, 100, 500, 1000).
@@
Discrete 2 IRMv1 49.47(0.21) 49.44(0.57)
4 49.77(0.78) 49.96(0.55)
8 49.60(0.65) 49.85(0.26)
16 49.74(0.62) 50.01(0.34)
100 49.62(0.58) 49.83(0.63)
500 50.28(0.87) 46.23(10.11)
1000 50.43(1.23) 49.63(15.06)
2 REx 82.95(0.47) 54.17(1.37)
4 81.93(0.91) 54.97(1.71)
8 81.59(0.89) 54.96(2.10)
16 81.60(0.69) 54.07(1.79)
100 80.94(0.78) 48.66(0.79)
500 81.48(0.67) 43.14(0.97)
1000 79.97(0.79) 42.24(0.74)
Table 7: Accuracy on the continuous CMNIST with $p_{s}(\bm{t})$ as a Sine function with different split number in (2, 4, 8, 16, 100, 500, 1000).
Results. Table 6 and 7 report the training and testing accuracy of different domain splitting schemes (2, 4, 8, 16) for both IRMv1 and REx on the continuous CMNIST dataset with two $p_{s}(\bm{t})$ settings. Under linear $p_{s}(\bm{t})$, REx performs the best with split number 8 by improving around 4% over the worst one. However, the results are fairly close under sine $p_{s}(\bm{t})$. On the other hand, IRMv1 with all testing accuracy around 0.5 does not seem to be able to extract useful features for the task with either linear $p_{s}(\bm{t})$ or sine $p_{s}(\bm{t})$.
@@
Discrete 2 IRMv1 72.68(0.36) 71.52(1.35)
5 73.46(1.41) 68.41(1.13)
10 72.97(1.88) 67.28(1.64)
25 74.42(2.10) 54.25(1.75)
50 75.40(0.99) 52.09(2.05)
2 REx 70.97(0.06) 74.43(1.32)
5 71.30(1.17) 74.42(0.87)
10 70.46(0.25) 73.20(1.65)
25 71.01(1.30) 72.96(0.24)
50 68.82(0.92) 72.90(0.46)
Continuous CIL 60.95(6.59) 76.25(4.99)
Table 8: Comparison on the real world dataset: HousePrice and Insurance. The metric is accuracy and standard deviations in parenthesis are calculated on three independent runs. Split Num stands for the number of domains we manually create by equally splitting the raw continuous domains.
@@
In this section, we tried different penalty set up in equation 3 on the auto-scaling dataset to validate the robustness of CIL. The approximating functions $h(\Phi(\bm{x})),g(\Phi(\bm{x}),\bm{y})$ are implemented by 2-Layer MLPs. We evaluate the performance changes by increasing either the hidden dimension of the MLPs (Table 9) or penalty weight $\lambda$ (Table 10).
Hidden Dimension ID Accuracy OOD Accuracy
32 84.37(5.95) 64.24(5.47)
64 84.17(2.58) 70.40(1.18)
128 81.25(1.65) 71.29(0.04)
256 85.57(2.67) 68.12(2.25)
512 85.28(2.27) 68.77(1.60)
Table 9: ID and OOD accuracy on the auto-scaling dataset across different MLP setups for $h(\Phi(\bm{x})),g(\Phi(\bm{x}),\bm{y})$ in equation 3. Standard deviation in brackets is calculated with 3 independent runs. Other settings are kept the same as in Section O
@@
100 86.91(1.31) 67.60(1.26)
1000 84.47(2.57) 70.01(1.54)
10000 84.17(2.58) 70.40(1.18)
100000 83.95(2.73) 70.42(1.16)
1000000 83.95(2.73) 70.35(1.24)
Table 10: ID and OOD accuracy on the auto-scaling dataset across different penalty weights $\lambda$ in equation 3. Standard deviation in brackets is calculated with 3 independent runs. Other settings are kept the same as in Section O
Results. The ID and OOD accuracy of different MLP setup (hidden dimension) for $h(\Phi(\bm{x})),g(\Phi(\bm{x}),\bm{y})$ under CIL on the auto-scaling dataset are shown in Table 9. CIL performs the best when the hidden dimension is 64, and the performance is stable with even higher dimensions. However, the testing accuracy drops to 64% when the dimension is reduced to 32, as the model is probably too simple to be able to extract enough invariant features. But it is still better than ERM shown in Table 3.
@@
  • LR: learning rate of the classification model $\Phi(\bm{x}))$, e.g. 1e-3.
  • OLR: learning rate of the penalty model $h(\Phi(\bm{x})),g(\Phi(\bm{x}),\bm{y})$, e.g. 0.001
@@
We show the parameter values used for each dataset in Table 11.